# Complements of discriminants of real parabolic function singularities. II

### 08 Mar 2026 v5

@@ 2.6. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the discriminant complement in some way. All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
@@
acts on the space $\Phi_{2}$. This operation always alters the real connected components of the discriminant complement, as well as the associated virtual components of type $\Phi_{2}$. Indeed, it maps any generic polynomial with value $I$ of the invariant Ind to a polynomial with value 3I-3-I. This operation is of order eight. All even iterations of it act as rotations by angles of $\frac{\pi k}{2}$ and thus belong to the group $G_{0}$. So, the entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation.
@@
Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [12], p. 160.
@@
The rest of the proof of Proposition 21 repeats the proof of Proposition 3 in § 5.4 of [28]. $\Box$
@@ 3. Components of discriminant complements for $J_{10}$ singularities
Proposition 26 (see [30]).
@@ 3.2. Components for $J_{10}^{1}$ singularities
No polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(f)$ can belong to the same connected component as its image f(x,y)f(x,-y) under the involution (33). Indeed, the polynomials (31) are of degree three in the variable $x$, so any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times. Therefore, the ovals placed on the different sides of the infinite component of the zero-level set cannot pass one above the other. On the other hand, the involution (34) moves the component of picture $(f)$ to the same component as (33).
The following is an invariant that separates each polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(e)$ from its mirror image f(x,y)f(x,-y). The four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take an arbitrary point inside each of these four ovals. There is a single curve with the equation $x=ay^{3}+by^{2}+cy+d$ that passes through these four points. The coefficient $a$ of this equation cannot be zero. Otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial in the coordinate $y$ of degree at most six with eight roots. Therefore, the sign of the coefficient $a$ does not depend on the choice of the four points inside the ovals and defines an invariant of the connected component of the discriminant complement. Any polynomial $f(x,y)$ in the considered component and its mirror image f(x,y)f(x,-y) define the opposite signs. The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
@@ 3.3. Components for $J_{10}^{3}$ singularities
Proof. The difference of all these 33 connected components follows from the topological considerations and Bezout’s theorem, see [25]. By Proposition 26, it remains to be shown that, in all cases not covered by the list (35), the involution (33) does not change the real connected component. The following polynomials (36), (37), (38), (39), are all invariant under this involution, so their virtual components are represented by single connected components of the complements of the discriminants.
@@
whose zero-level set is shown in Fig. 6. Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal weighted homogeneous part will be of class $J_{10}^{3}$. The obtained polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values. It will also have two critical points in the complex domain. Then we add a very small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain. The zero-level set of the resulting polynomial will have the desired shape shown in picture $(e)$ of Fig. 5, see Fig. 6 right. Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal weighted homogeneous part of the resulting polynomial will have the canonical form, see (32). Finally, we apply a parallel translation of the $x$ coordinate to kill the monomial $x^{2}$. The resulting polynomial realizes the picture $(e)$. It is invariant under the involution (33), in particular it has the form (32) with $\lambda_{2}=\lambda_{4}=\lambda_{7}=\lambda_{9}=0$. $\Box$
@@ 4. The case of $P_{8}$ singularities
Two classes $P_{8}^{1}$ and $P_{8}^{2}$ of function singularities in three variables have the canonical normal form
@@
with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$, see [5]. The zero-level sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected component in ${\mathbb{R}P}^{2}$. Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to a plane and has a single singular point at the origin. The zero-level set of a polynomial of class $P_{8}^{2}$ consists of a surface like this and a surface that is homeomorphic to the standard quadratic cone.
@@
A polynomial $f:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all distinct and not equal to 0, and the complex zero-level set of the principal homogeneous part of $f$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
@@
It follows easily from Morse theory that the invariant Ind can take only the values 1,0,1,2,1,0,-1,-2, and 3-3 on polynomials of classes $\Xi_{1}$ and $\Xi_{2}$. The closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f_{\lambda}$ of a singularity (41) with $\mbox{\rm Ind}(f_{\lambda})$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, three handles, or the disjoint union of the projective plane and the sphere)).
@@ 4.2. The homological invariant of real components
Clearly, the invariant Ind is equal to the alternating sum of three elements of the homology index.
@@ 4.3. The symmetry group and the relation between the real and virtual components
The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following alterations: the frames of vanishing cycles in the groups $H_{2}(f^{-1}(0))$ for non-discriminant $f\in\Xi_{1}$ or $\Xi_{2}$ consist of eight elements, and the data sets of virtual functions only include the parities of the Morse indices of critical points, and not their integer values. We also have the following analog of Proposition 3.
@@
The group $S(3)$ of coordinate permutations in ${\mathbb{R}}^{3}$ preserves the normal form (41) and acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. This action is generated by the following transformations:
@@
In particular, this group can act on the set of the connected components of the discriminant complements in the spaces $\Xi_{i}$. The transformations (44) preserve all topological characteristics of the functions. Therefore, all these connected components within the same orbit of this action are associated with the same virtual component of the formal graph of type $P_{8}^{1}$ or $P_{8}^{2}$.
@@
Proof. According to Lemma 38 of [29], there are only six polynomials of the form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$. Three of these polynomials (mentioned in formula (27) of [29] with a certain value of the parameter $\beta$) have signature (0,2) at the $A_{2}$ point. The remaining three polynomials have the form (46) up to the permutations of coordinates, see formula (28) in [29]. In particular, they belong to one orbit of the group $S(3)$. $\Box$
@@
The set of all polynomials of of the form (42) and class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
@@
Proof of Proposition 35. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 36 (respectively, 37) repeats the deduction of Proposition 9 from Lemma 10. Specifically, the neighborhood of the polynomial (46) or (47) in $\Xi_{2}$ has only one non-trivial automorphism that commutes with the Lyashko–Looijenga map: the involution $f(x,y,z)\leftrightarrow f(y,x,z)$. $\Box$
@@ 4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
The two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2. For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$. According to the last sentence of the previous paragraph, the associated virtual components are also different.
@@
have $\mbox{Ind}=-3$. They are related to each other by the involution (45). Therefore, according to the final statement of Proposition 38, they realize two distinct virtual components. Since these polynomials are invariant under the entire group $S(3)$, these virtual components are also associated with single real components. Thus, Theorem 39 is proved for all virtual components with Ind equal to 1,0,2,1,0,-2, and 3-3. Let us prove it for the only remaining virtual component with $\mbox{Ind}=-1$.
@@
Denote by $C(l)$ the critical value of the polynomial (52) at its Morse critical point. Let $F(l)$ be the polynomial obtained from (52) by adding the constant $C(l)/2$. Then the polynomial $F(l)$ is non-discriminant and has $\mbox{Ind}=-1$.
The latter polynomial is canonically determined by the linear form $l$ and depends continuously on it. The union of polynomials $F(l)$ over all $l\in L$ is connected and invariant under the action of the group $S(3)$. Therefore, the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement. $\Box$
@@ 4.5. One-dimensional cohomology of the discriminant complement of $P_{8}^{1}$ singularities
Thus, the one-dimensional homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$. These homology groups form a local system over the space of all such polynomials, i.e., over the connected component of the discriminant complement under consideration. In particular, the fundamental group of this component acts on these groups by monodromy. For each $l\in L$, one basic homology class of the zero-level set of the polynomial $F_{l}$ is defined by the intersection of this set and the plane $\ker l$. When $l\in L$ runs a half of the circle $L$, the planes $\ker l$ run a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomials $F(l)$ run a closed cycle in the discriminant complement. The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation. Thus, the half-circle $L/{\mathbb{Z}}_{2}$ defines non-trivial elements of the fundamental group and of the mod 2 homology group of the component. $\Box$

### 16 Mar 2026 v6

@@ 2.6. Components for the $X_{9}^{2}$ singularities
The group $G_{0}$ (see § 2.3) acts on the space $\Phi_{2}$ of all polynomials (1) with $f_{A}=x^{4}+2Ax^{2}y^{2}+y^{4},$ A<1A<-1, preserving the coefficient $A$ and permuting the connected components of the discriminant complement. All these components belonging to the same orbit of this action are associated with the same virtual component of the formal graph of type $X_{9}^{2}$.
@@
acts on the space $\Phi_{2}$. This operation always alters the real connected components of the discriminant complement, as well as the associated virtual components of type $\Phi_{2}$. Indeed, it maps any generic polynomial with value $I$ of the invariant Ind to a polynomial with value 3I-3-I. This operation is of order eight. All even iterations of it act as rotations by angles of $\frac{\pi k}{2}$ and thus belong to the group $G_{0}$. The entire set of connected components of the discriminant complement in the space $\Phi_{2}$ splits into the orbits of this operation.
@@
Proof. This follows from the well-known formula for $j$-invariant of quartic forms, see e.g. [18]. $\Box$
@@
The rest of the proof of Proposition 21 repeats the proof of Proposition 9 with the simplification indicated in the proof of Proposition 16. Let $f_{0}$ be one of eight polynomials mentioned in in Lemma 22. The space $\Phi_{2}$ does not have automorphisms that commute with the Lyashko–Looijenga map and map this polynomial to itself, see the proof of Proposition 3 in § 5.4 of [29]. Let $f$ and $\tilde{f}$ be two generic polynomials of class $\Phi_{2}$ with the same associated virtual function and sets of critical values and paths. Let $I$ be a generic path in $\Phi_{2}$ connecting $f$ with $f_{0}$, and $\tilde{I}$ be a path starting at $\tilde{f}$ and repeating all the critical values and surgeries of the path $I$. Then the endpoint $\tilde{f}_{0}$ of this path is well-defined and is also one of these eight points, and the entire path $\tilde{I}$ is mapped to $I$ by the element of the group $G_{0}$ that maps the point $\tilde{f}_{0}$ to $f_{0}$. $\Box$
@@ 3. Components of discriminant complements for $J_{10}$ singularities
The following analog of Propositions 2 and 3 follows from the same considerations as these statements.
Proposition 26.
For any class of polynomials $\Omega_{1}$ or $\Omega_{3}$, 1)) if two generic polynomials, $f$ and $\tilde{f}$, of this class belong to the same connected component of the complement of the discriminant, then the virtual components $S(f)$ and S(f~)S($\tilde${f}) are the same.
2)) every virtual function of this class is associated with a generic real polynomial of the same class;
3)) for any generic polynomial $f$ of this class, every virtual function from the virtual component $S(f)$ is associated with some generic real polynomial $\tilde{f}$ from the same connected component of the complement of the discriminant as $f$. $\Box$
Proposition 27 (see [31], Proposition 9).
@@ 3.2. Components for $J_{10}^{1}$ singularities
No polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(f)$ can belong to the same connected component as its image f(x,y)f(x,-y) under the involution (33). Indeed, the polynomials (31) are of degree three in the variable $x$, so any vertical line $\{y=\mbox{const}\}$ can intersect the curve $\{f=0\}$ at most three times. Therefore, the ovals placed on the different sides of the infinite component of the zero-level set cannot pass one above the other. On the other hand, the involution (34) moves the component of picture $(f)$ to the same component as the involution (33).
Every polynomial $f(x,y)$ of class $\Omega_{1}$ that realizes the picture $(e)$ belongs to a different connected component than its mirror image f(x,y)f(x,-y). Indeed, the four ovals defined by this polynomial are canonically ordered from the right to the left, and this ordering is continuous along the connected components of the discriminant complement. Let us take an arbitrary point inside each of these four ovals. There is a single curve with the equation $x=ay^{3}+by^{2}+cy+d$ that passes through these four points. The coefficient $a$ of this equation cannot be zero. Otherwise, the restriction of the polynomial $f$ to this curve would be a polynomial in the coordinate $y$ of degree at most six with at least eight roots. Therefore, the sign of the coefficient $a$ does not depend on the choice of the four points inside the ovals and remains constant along the connected component of the discriminant complement. Any polynomial $f(x,y)$ in the considered component and its mirror image f(x,y)f(x,-y) define opposite signs. The same is true for polynomials that realize the picture (e~)($\tilde${e}). $\Box$
@@ 3.3. Components for $J_{10}^{3}$ singularities
Proof. The difference of all these 33 connected components follows from the topological considerations and Bezout’s theorem, see [26]. By Proposition 27, it remains to be shown that, in all cases not covered by the list (35), the involution (33) does not change the real connected component.
The following polynomials (36), (37), (38), (39), are all invariant under this involution, so their virtual components are represented by single connected components of the complements of the discriminants.
@@
whose zero-level set is shown in Fig. 6. Then, we add the term $\varepsilon y^{6}$ with sufficiently small $\varepsilon>0$ so that the principal weighted homogeneous part will be of class $J_{10}^{3}$. The obtained polynomial will then keep eight real critical points (three minima and five saddlepoints), which are visible in this figure, and do not significantly change their critical values. It will also have two critical points in the complex domain. Then we add a small positive constant to the obtained polynomial in such a way that all saddlepoints will have positive critical values and all minima will stay in the negative domain. The zero-level set of the resulting polynomial will have the desired shape shown in picture $(e)$ of Fig. 5, see Fig. 6 right. Then we apply a coordinate change of the form $\tilde{x}=x+\alpha y^{2},\tilde{y}=\beta y$ in such a way that the principal weighted homogeneous part of the resulting polynomial will have the canonical form, see (32). Finally, we apply a parallel translation of the $x$ coordinate to kill the monomial $x^{2}$. The resulting polynomial realizes the picture $(e)$. It is invariant under the involution (33), in particular it has the form (32) with $\lambda_{2}=\lambda_{4}=\lambda_{7}=\lambda_{9}=0$. $\Box$
@@ 4. The case of $P_{8}$ singularities
The two classes $P_{8}^{1}$ and $P_{8}^{2}$ of function singularities in three variables have the canonical normal form
@@
with $A<1$ for $P_{8}^{1}$ singularities and $A>1$ for $P_{8}^{2}$, see [5]. The zero-level sets of these normal forms of class $P_{8}^{1}$ (respectively, $P_{8}^{2}$) define cubic curves with one (respectively, two) connected component in the projective plane. In the case of two components, one of them is a convex oval. The point $(1:1:1)$ lies in the disc bounded by this oval, because both the curve and this point are invariant under the action of the group $S(3)$ of coordinate permutations in ${\mathbb{R}}^{3}$. Thus, each polynomial of class $P_{8}^{1}$ defines a conical surface in ${\mathbb{R}}^{3}$ that is homeomorphic to a plane and has a single singular point at the origin. The zero-level set of a polynomial of class $P_{8}^{2}$ consists of a surface like this one and a surface that is homeomorphic to the standard quadratic cone.
@@
A polynomial $f_{\lambda,A}:({\mathbb{C}}^{3},{\mathbb{R}}^{3})\to({\mathbb{C}},{\mathbb{R}})$ of the form (42) is generic if it has only Morse critical points in ${\mathbb{C}}^{3}$, its critical values are all distinct and not equal to 0, and the complex zero-level set of its principal homogeneous part $f_{A}$ defines a smooth curve in ${\mathbb{C}}P^{2}$.
@@ 4.2. The homological invariant of real components
Clearly, the homology index is an invariant of the connected components of discriminant complement, and the invariant Ind is equal to the alternating sum of three elements of the homology index.
The values that the invariant Ind can take on polynomials of classes $\Xi_{1}$ and $\Xi_{2}$ are described in Propositions 39 and 43 below. These values are 1,0,1,2,1,0,-1,-2, and 3-3. The closure in ${\mathbb{R}P}^{3}$ of the set of zeros of any non-discriminant perturbation $f$ of a singularity (41) with $\mbox{\rm Ind}(f)$ equal to 0 ((respectively, 1-1, 2-2, 3-3, or 1)) is homeomorphic to the projective plane ((respectively, the projective plane with one, two, or three handles, or the disjoint union of the projective plane and the sphere)).
@@ 4.3. The symmetry group and the relation between the real and virtual components
The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following alterations: the frames of vanishing cycles in the groups $H_{2}(f^{-1}(0))$ for non-discriminant polynomials $f\in\Xi_{1}$ or $\Xi_{2}$ consist of eight elements, and the data sets of virtual functions only include the parities of the Morse indices of critical points, and not their integer values. We also have the following analog of Proposition 3.
@@
The group $S(3)$ of coordinate permutations in ${\mathbb{R}}^{3}$ acts on the spaces $\Xi_{1}$ and $\Xi_{2}$ preserving the normal form (41) and the discriminant variety. This action is generated by the following transformations:
@@
Consequently, this group acts on the set of the connected components of the discriminant complements in the spaces $\Xi_{*}$. The transformations (44) preserve all topological characteristics of the functions. Therefore, all these connected components within the same orbit of this action are associated with the same virtual component of the formal graph of type $P_{8}^{1}$ or $P_{8}^{2}$.
@@
Proof. According to Lemma 38 of [30], there are only six polynomials of the form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$. Three of these polynomials (mentioned in formula (27) of [30] with a certain value of the parameter $\beta$) have signature (0,2) of the second differential at the $A_{2}$ point. The remaining three polynomials have the form (46) up to the permutations of coordinates, see formula (28) in [30]. In particular, they belong to one orbit of the group $S(3)$. $\Box$
@@
The set of all polynomials of of the form ((42)) and class $\Xi_{2}$ that have a critical point of class $E_{6}$ with zero critical value, and also two real critical points of class $A_{1}$ with the same critical value $\frac{64}{3}+\frac{128}{3\sqrt{3}}$, consists of three polynomials: the polynomial
@@
Proof of Proposition 36. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 37 (respectively, 38) repeats the deduction of Proposition 9 from Lemma 10. Specifically, the neighborhood of the polynomial (46) or (47) in $\Xi_{2}$ has only one non-trivial automorphism that commutes with the Lyashko–Looijenga map: the involution $f(x,y,z)\leftrightarrow f(y,x,z)$. $\Box$
@@ 4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
The two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2. For the first component, this invariant is equal to $(1,0,0)$, and for the second it is equal to $(0,0,1)$. According to the last sentence of the previous paragraph, the associated virtual components are also different.
@@
have $\mbox{Ind}=-3$. These polynomials are invariant under the entire group $S(3)$. Therefore, the virtual components associated with them are associated only with the real components of discriminant complements that contain these polynomials. These two polynomials are related to each other by the involution (45). Therefore, according to the final statement of Proposition 39, their associated virtual components are different. Thus, Theorem 40 is proved for all virtual components with Ind equal to 1,0,2,1,0,-2, and 3-3. Let us prove it for the only remaining virtual component with $\mbox{Ind}=-1$.
@@
Denote by $C(l)$ the critical value of the polynomial (52) at its Morse critical point. Let $F(l)$ be the polynomial obtained from (52) by subtracting the constant $C(l)/2$. Then the polynomial $F(l)$ is non-discriminant and has $\mbox{Ind}=-1$.
The latter polynomial is canonically determined by the linear form $l$ and depends continuously on it. The union of polynomials $F(l)$ over all $l\in L$ is connected and invariant under the action of the group $S(3)$. Therefore, the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement. This completes the proof of Theorem 40. $\Box$
@@ 4.5. One-dimensional cohomology of the discriminant complement of $P_{8}^{1}$ singularities
Thus, the one-dimensional homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$. These homology groups form a local system over the space of all such polynomials, i.e., over the connected component of the discriminant complement under consideration. In particular, the fundamental group of this component acts on these groups by monodromy. For each $l\in L$, one basic homology class of the zero-level set of the polynomial $F(l)$ is defined by the intersection of this set and the plane $\ker l$. When $l\in L$ runs a half of the circle $L$, the planes $\ker l$ run a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomials $F(l)$ run a closed cycle in the discriminant complement. The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation. Thus, the half-circle $L/{\mathbb{Z}}_{2}$ defines non-trivial elements of the fundamental group and of the mod 2 homology group of the component. $\Box$